A U-tube manometer has Mercury of relative density off 13.68 as its manometric liquid. Kerosene of relative density of 0.6 was poured into one of the arms and so a drop of 2.0 cm results in that arm. Also, water with a relative density of 1g/cm3 was poured into the other arm and so the two arms became the same again. Calculate 1 the height of the kerosene to the height of the water column in details.
1
Expert's answer
2021-07-16T10:19:48-0400

As per the given question,

Density of mercury "(\rho)=13.68"

Relative density of kerosene "(\rho_k)=0.6"

Drop in height "(\Delta H)= 2 cm"

Relative density of water "(\rho_w)=1g/cm^3"

Height of kerosene =?

Height of water = ?

"\rho_1 h_1=\rho_2h_2"

Now, substituting the values,

"h_1=\frac{\rho_2 h_2}{\rho_1}"


"=\frac{0.6\times 2}{13.68}cm"

"=0.0877cm"